Im am puzzled with the steps I need to solve, can anyone help?
Determine (a) the area and (b) the circumference of the circle
circle diameter is 28yd.

Answer by rfer(16322)

A=pir^2
A=3.14*14^2
A=615.44
-------------
C=pid
C=3.14*28
C=87.92